2018 Oct 17
Measurements on-sky show an added rotation of ~1 degree for the f/32 camera due to the mirror shift, when measuring the instrument alignment angle during on-sky checkouts afte re-installation. We have updated the telescope Instrument Alingment Angle (IAA) accordingly. The IAA is used by the telescope control system to position the cass rotator and thus the physical rotation of the instrument relative to sky to achieve the correct instrument position angle that is requested for each observation. With the adjustment to the IAA, there should be minimal impact on science data, as we can measure the IAA to the same precision as previously. Due to this we are also in the process of remeasuring the Altair field lens barrel distortion mapping to provide a new solution if one is needed.
2018 Oct 2
We encountered an issue with one of the computer controller cards that reads out NIRI images and sends the data to the Gemini servers on September 6. NIRI was not able to read out images at that time. In mid-September,we were able to recover the instrument, and NIRI is taking images again.The root cause of the fault was determined to be a voltage issue with the detector controller electronics. At the end of the planned September Gemini North Engineering Shutdown, NIRI was installed on the telescope. We find after the August instrument engineering work, that the alignment of the f/32 camera has shifted slightly, but we are able to compensate for this with the telescope pointing. At this time, we do not find any impact to science observations. We will continue to characterize NIRI performance and the f/32 camera, and we will update the NIRI documentation and inform program PIs as needed
2018 Aug 2
The pupil viewer that we use to help align the instrument on the telescope, is not consistently and reliably coming into position. Engineering work will be performed ths month to improve the pupil viewer performance.
2018 May 10
NIRI overheads were recently measured, and the new values have been updated on the public website.
2012 Feb 25
NIRI is available for queue and classical imaging observations.
Not all filters are available at any one time; check the filter list to see which are currently installed.
NIRI occasionally suffers from vertical striping and shifts in DC levels which can both vary from quadrant to quadrant. Details can be found here.
2010 Jun 21
The NIRI focal plane unit which holds the various spectroscopic slits and imaging masks stopped responding on the night of June 19. On Sunday we were able to nudge the wheel to the f/6 imaging position and today engineers locked out the mechanism so that NIRI may be safely used for imaging with any camera in combination with the f/6 beam splitter and f/6 focal plane mask. Spectroscopy will be unavailable until NIRI can be warmed up and taken into the lab for repairs. Until we can disassemble NIRI to assess the severity of the problem it is unknown how much time will be required to return NIRI to a fully operational state.
2010 Apr 29
The beam splitter wheel, which determines the size of the beam to direct to the science detector, is again having problems moving into position reliably. In order to resume science operations with NIRI, we have decided to fix the beam splitter at the f/6 position which sends the largest field of view through the instrument and can therefore be used with all three cameras. This results in a several percent increase in background (which is not uniform over the field) and slightly different vignetting when used with Altair and its field lens. See here for more details.
2010 Apr 15
NIRI is cold and the beam splitter mechanism is working perfectly! NIRI is about to be re-installed on the telescope and should be available for science tomorrow.
2010 Apr 2
The beam splitter wheel detent roller and bushing were removed and refurbished (the inner bushing had translated axially and was binding the roller against the detent arm) and the Phytron drive motor was replaced. While NIRI was open we recut the f/32 4-pixel slit to be parallel and vertical.
2010 Mar 24
NIRI is cold again, and there are now no slipping gears, however, there is still too much friction in the system for the motor to move the beam splitter wheel reliably while cold. The two likely culprits which could be causing this problem are the Geneva detent mechanism roller and/or the drive motor. We have therefore begun another warm up in order to work on these components. If all goes well NIRI should be available for science in ~3 weeks. More information will be posted here as it becomes available.
2010 Mar 15
The beam splitter wheel gear box was removed and disassembled. An inspection uncovered two places where the gears had spun on their shafts, so new set-screws were installed, including one which was bent and rubbing against a nearby gear. The Geneva gear and roller assembly was re-lubricated and the gear box was re-installed. We are continuing with a thorough examination of other mechanisms while NIRI is in the lab, and will start cooling the dewar on Thursday March 18, with an on-sky target date of March 24.
2010 Mar 4
The NIRI beam-splitter wheel which is used to select between the three cameras (optomechanical layout) stopped responding on the afternoon of March 2. Troubleshooting indicates that the problem is internal and mechanical. We will therefore begin warming NIRI today to investigate further and expect NIRI to be unavailable for science through the end of March.
2009 Sep 10
After a very thorough investigation, Gemini engineers discovered two problems with NIRI: the detector controller electronics had an intermittent issue with the slow clocks, and there was a short circuit between two pins in the detector mount. These problems were fixed and routine maintenance was performed on the detector mount and cabling. Testing is ongoing, and NIRI is now being scheduled for science operations.
2009 Aug 27
Work on the NIRI detector and electronics is ongoing, and we are planning on cooling down next week to do more testing. The earliest NIRI may return to science operations is September 10.
2009 Aug 12
Over the past four months we have seen several failures of the NIRI detector to activate. These events lasted anywhere from several minutes to several hours, and the instrument was unusable during these periods. We contacted outside experts and were investigating options for troubleshooting the problem.
Unfortunately, on the night of 2009 Aug 9 this condition transitioned from intermittent to persistent, and we have not been able to activate the array since.
As the cause and severity of the problem are still unknown, it is impossible to provide an exact date when NIRI will be fixed. We are starting to warm up NIRI today so that we may perform diagnostics on the detector and control electronics. We are hoping that the problem can be located and resolved quickly so that NIRI can return to science operations by the end of August. However, NIRI may be unavailable for longer if invasive work is required.
2009 July 14
Filters rearranged for 2009B. Removed H2O ice, hydrocarbon, Br(alpha)cont. Installed Hcont, CH4(long) and HeI(2p2s).
2009 Jan 15
Filters rearranged for 2009A. Remove Bra, CH4long, Hcon157. Installed H2Oice, hydrocarb, Bracont.
2008 July 25
Filters rearranged for 2008B. Removed Jcont(1.12), Pa(beta), K(short) and hydrocarbon. Installed Y, Br(alpha), and two NB ice filters.
2008 May 14
Changed L spectroscopic flat configurations in OT library to give narrower range of signals across band.
2006 May 22
Replaced cold heads.
2005 Dec 12
Removed FeII filter; installed Pabeta and Jcont(1207).
2005 May 9
Removed Jcont(1.207) and Pabeta filters; installed Jcont(1.065).
Begin multi-instrument queue observing.
2004 Nov 5
The Oct 21 - Nov 4 run was nearly wiped out by bad weather, with clouds and high humidity during the entire run. The problems with the array readout culminated during this run, and on Oct 25 we changed out the coadder board, and on Nov 1 we disabled the window cover ionizer. Ever since the array has been behaving very well, and we have seen very few instances of the quadrant bias variations or pattern noise.
2004 May 16
Third and fourth NIRI runs completed. Weather generally better during both runs, but with only a few nights of excellent seeing. ALTAIR now in frequent use with NIRI, with almost all Band 1-3 programs under way or completed. Problems with electronic pickup on array frames were somewhat alleviated starting on April 5. Typically 90% of the data frames obtained in each program since then are uncontaminated; we continue to work toward a complete solution to this problem. One of NIRI's internal motors began failing toward the end of the May run, limiting our observing options during the final three nights of the May run. That motor is being replaced during mid-May and it is expected that the instrument will be available for use at the start of the May 29 - June 9 run.
2004 Mar 12
Second 2004A observing run completed - weather conditions were appallingly bad - heavy snow/thunderstorm complete with electrical damage closed telescope for several nights; weather poor thereafter (clouds, horrible seeing) - only able to do parts of the few programs that could tolerate those conditions. Continued problems with vertical striping and different dc levels in the four array quadrants in some images; as the pattern appears nearly constant, some workarounds for this are possible.
2004 Feb 12
First 2004A observing run completed - weather conditions were generally poor (esp seeing). Some problems with array dc level instability visible as vertical striping and as different dc levels in the four array quadrants in some images; otherwise instrument performance was nominal.
2004 Jan 20
NIRI update, Dewar being cooled after mechanical overhaul and replacement of some filters in preparation for 2004B
2003 Jun 19
NIRI update, Semester 2003A Results
2003 Jan 31
NIRI update, Important guidance on selecting guide stars for NIRI during 2003A
2002 Dec 29
NIRI completes scheduled queue nights for 2002B - weather (seeing) not great on most December nights, but instrument performs well
2002 Dec 2
NIRI on telescope and being used in H-band for commissioning of ALTAIR
2002 Oct 7
NIRI update, damaged mechanisms inspected and fixed
2002 Sep 24
NIRI update; failure of pupil wheel and gimbal mechanism
2002 Sep 12
NIRI update, beginning of semester 2002B
2002 Jul 12
NIRI update, end of semester 2002A
2002 Apr 10
NIRI update, midway through semester 2002A
2002 Jan 30
Report on progress with NIRI during semester 2001B; update on current status
2001 Sept 26
Excellent progress with NIRI flexure; update on acceptance tests, queue and System Verification observations
2001 May 3
Statement regarding the availability of NIRI in 2001A
News Archive Filter
The GEMMA Podcast
A podcast about Gemini Observatory and its role in the Era of Multi-Messenger Astronomy. Featuring news related to multi-messenger astronomy (MMA), time-domain astronomy (TDA), our visiting instrument program, and more through interviews with astronomers, engineers, and staff both here at Gemini (North and South) and abroad.